OpenStudy (anonymous):

What is the vertex of y = |x + 1| ? (–1, 0) (0, –1) (0, 1) (1, 0)

OpenStudy (anonymous):

not completely. I thought it was D.

OpenStudy (anonymous):

x and 1 ? O.o

hartnn (hartnn):

ok, let me make it easier y= |x-h| +k vertex = (h,k)

hartnn (hartnn):

here u have |x+1| so h=-1 got this?

hartnn (hartnn):

and you have nothing added to |x+1| so k=0 ok?

OpenStudy (anonymous):

okay... lol

hartnn (hartnn):

so vertex is (h,k) = (-1,0) did u understand how?

OpenStudy (anonymous):

not at all and i have 5 more problems :(

hartnn (hartnn):

 now ?